# DriveMind: A Dual-VLM based Reinforcement Learning Framework for Autonomous Driving

### v1

@@ Abstract
End-to-end autonomous driving systems map sensor data directly to control commands, but remain opaque, lack interpretability, and offer no formal safety guarantees. While recent vision-language-guided reinforcement learning (RL) methods introduce semantic feedback, they often rely on static prompts and fixed objectives, limiting adaptability to dynamic driving scenes. We present DriveMind, a unified semantic reward framework that integrates: (i) a contrastive Vision-Language Model (VLM) encoder for stepwise semantic anchoring; (ii) a novelty-triggered VLM encoder-decoder, fine-tuned via chain-of-thought (CoT) distillation, for dynamic prompt generation upon semantic drift; (iii) a hierarchical safety module enforcing kinematic constraints (e.g., speed, lane centering, stability); and (iv) a compact predictive world model to reward alignment with anticipated ideal states. DriveMind achieves 19.4$\pm$2.3 km/h average speed, 0.98$\pm$0.03 route completion, and near-zero collisions in CARLA Town 2, outperforming baselines by over 4% in success rate. Its semantic reward generalizes zero-shot to real dash-cam data with minimal distributional shift, demonstrating robust cross-domain alignment and potential for real-world deployment.
@@ 1 Introduction
We introduce DriveMind, which comprises four key modules, each targeting a key limitation of existing approaches. First, a contrastive VLM provides dense, context-aware semantic rewards by anchoring each bird’s-eye-view (BEV) frame against fixed “present” and “ideal” concepts, addressing the lack of informative feedback and poor context adaptation in traditional reward designs. Second, a novelty-triggered VLM encoder–decoder is invoked only when embedding drift exceeds a threshold, generating adaptive “present” (hazard) and “ideal” (goal) prompts to prevent reward hacking in repetitive or adversarial scenarios. This on-demand prompting with asynchronous chain-of-thought (CoT) distillation using GPT-4 minimizes per-step computation, refreshing embeddings only upon genuine scene changes, hence having negligible impact on the overall latency while maintaining efficiency. Third, a hierarchical fusion module multiplicatively combines normalized kinematic factors, such as speed regulation, lane centering, heading alignment, lateral stability, to enforce a hard safety veto whenever any physical constraint is violated. Fourth, a compact world model forecasts the next semantic embedding and provides a predictive contrastive foresight reward, improving long-horizon credit assignment and anticipatory planning. Together, these modules yield an adaptive, interpretable, and provably safer reward framework for autonomous driving. Experiments on CARLA Town 2 demonstrate improved route completion, collision avoidance, and sample efficiency. Zero-shot transfer to real dash-cam footage shows robust cross-domain alignment, confirming DriveMind’s promise for safer, more interpretable end-to-end driving.
@@
  • We validate DriveMind via extensive experiments on CARLA Town 2 and zero-shot transfer to BDD100K dash-cam data, achieving $19.4\pm 2.3$ km/h average speed, $0.98\pm 0.03$ route completion, near-zero collision speed in simulation, and minimal distributional shift in real data (Wasserstein $=0.028$, KS $=0.105$), surpassing baselines on most safety and efficiency metrics.
@@ 3 Proposed Approach: DriveMind
Figure 1: Dual-VLM architecture and reward pipeline of DriveMind: A static contrastive VLM encoder maps each bird’s-eye-view image $\psi(s_{t})$ to a fixed semantic embedding $v_{t}$, while a novelty detector asynchronously triggers a dynamic VLM encoder-decoder to update “present” and “ideal” prompt embeddings $(u^{P}_{t},u^{I}_{t})$ via distilled GPT-4 chain-of-thought. These embeddings feed into the Reward Combiner, yielding (i) Adaptive Ideal-State Contrastive Reward, (ii) Hierarchical Vehicle-State Fusion Reward, and (iii) Predictive Contrastive Foresight Reward.
@@ 3.4.3 Predictive Contrastive Foresight Module
Combining all components yields the final per-step reward:
@@ 4.1 Experimental Setup
Observations & Actions. At each timestep, the agent receives an observation $o_{t}$ consisting of three parts. First, a bird’s-eye-view semantic image, resized to $224\times 224$ px, fuses camera and LiDAR segmentation of lanes, vehicles, pedestrians, and obstacles. Second, an ego-state vector $e_{t}=(v_{t},\theta_{t},\omega_{t})$ captures normalized longitudinal speed, steering angle, and yaw rate from onboard sensors. Third, a sequence of $K=10$ future waypoints $(x_{t+1},y_{t+1},\dots,x_{t+K},y_{t+K})$, sampled every 5 m in the vehicle’s local frame and scaled to $[-1,1]$, provides route guidance. Together, these inputs form the full policy observation. The agent outputs a continuous action $a_{t}=(a^{(1)},a^{(2)})\in[-1,1]^{2}$, where $a^{(1)}$ maps linearly to steering angle and $a^{(2)}$ splits into throttle (positive) and brake (negative), each normalized to $[0,1]$. This compact interface enables real-time, smooth control in response to learned reward signals.
@@
Category Model AS $\uparrow$ RC $\uparrow$ TD $\uparrow$ CS $\downarrow$ SR $\uparrow$ Expert-Designed (Binary) TIRL-SAC [26] 0.37$\pm$0.28 0.01$\pm$0.001 4.7$\pm$3.5 0.24$\pm$0.34 0.00$\pm$0.00 TIRL-PPO [26] 0.43$\pm$0.23 0.01$\pm$0.005 14.8$\pm$9.7 0.10$\pm$0.15 0.00$\pm$0.00 Expert-Designed (Summation) Chen-SAC [27] 21.4$\pm$1.16 0.29$\pm$0.12 663.6$\pm$286.7 2.07$\pm$2.21 0.08$\pm$0.08 ASAP-RL-PPO [15] 1.25$\pm$0.30 0.01$\pm$0.00 28.1$\pm$3.4 0.61$\pm$0.61 0.00$\pm$0.00 ChatScene-SAC [28] 17.7$\pm$0.12 0.88$\pm$0.03 1763.2$\pm$90.9 1.18$\pm$0.46 0.73$\pm$0.05 ChatScene-PPO [28] 15.3$\pm$0.33 0.78$\pm$0.05 1515.6$\pm$129.1 0.89$\pm$0.32 0.63$\pm$0.05 LLM-Based Revolve [29] 18.4$\pm$0.03 0.92$\pm$0.11 1915.3$\pm$248.3 1.53$\pm$2.16 0.83$\pm$0.24 Revolve-Auto [29] 17.2$\pm$0.76 0.80$\pm$0.06 1539.6$\pm$147.5 1.65$\pm$0.28 0.63$\pm$0.05 VLM-Based VLM-SR [7] 0.53$\pm$0.27 0.02$\pm$0.00 47.9$\pm$9.2 0.18$\pm$0.25 0.00$\pm$0.00 RoboCLIP [9] 0.44$\pm$0.05 0.07$\pm$0.03 146.3$\pm$62.3 1.05$\pm$0.58 0.00$\pm$0.00 VLM-RM [8] 0.20$\pm$0.05 0.02$\pm$0.01 35.9$\pm$25.8 0.003$\pm$0.005 0.00$\pm$0.00 LORD [19] 0.17$\pm$0.08 0.02$\pm$0.02 45.1$\pm$57.1 0.02$\pm$0.02 0.00$\pm$0.00 LORD-Speed [19] 18.9$\pm$0.36 0.87$\pm$0.05 1783.4$\pm$172.8 2.80$\pm$1.16 0.67$\pm$0.05 VLM-RL [20] 19.3$\pm$1.29 0.97$\pm$0.03 2028.2$\pm$96.60 0.02$\pm$0.03 0.93$\pm$0.04 Ours DriveMind 19.4$\pm$2.34 0.98$\pm$0.03 2083.1$\pm$50.09 0.01$\pm$0.07 0.97$\pm$0.06
@@ 4.3 Ablation Studies
DriveMind, NoHVFR $0.58\pm 0.25$ $0.02\pm 0.00$ $0.14\pm 0.21$ $00.00\pm 0$ 0.00
DriveMind, NoPCFM $19.2\pm 1.14$ $0.92\pm 0.05$ $0.02\pm 0.02$ $0.87\pm 0.06$
@@ 4.4 Real-World Generalization
Table 3 compares the Adaptive Ideal-State Contrastive Reward (AICR) distributions in CARLA Town 2 versus 10,000 BDD100K bird’s-eye-view frames. On real data, the mean reward drops slightly from 0.130 to 0.118 and the standard deviation rises modestly from 0.080 to 0.085. The fraction of positive-reward frames falls from 88.0 % to 77.4 %, while negative-reward frames increase from 12.0 % to 22.6 %. Crucially, the Wasserstein-1 distance (0.028) and the Kolmogorov-Smirnov statistic (0.105) remain low, indicating that the two empirical distributions closely align. These results confirm that DriveMind’s learned semantic objectives transfer effectively, maintaining robust AICR behavior in unconstrained, real-world driving scenes without any additional fine-tuning.
@@
Dataset Mean $r_{\rm AICR}$ STD $r_{\rm AICR}$ $r_{\rm AICR}>0$ $r_{\rm AICR}<0$ EMD (CARLA $\rightarrow$ Real) KS-stat CARLA Town 2 0.130 0.080 88.0% 12.0% , , BDD100K (10K frames) 0.118 0.085 77.4% 22.6% 0.028 0.105
@@ Appendix A Preliminaries
A.1 Partially-Observable Markov Decision Processes and Soft Actor-Critic
We model autonomous driving as a partially-observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},R,\gamma)$, where $\mathcal{S}$ is the hidden state space, $\mathcal{A}$ the continuous control actions, and $\mathcal{O}$ the agent’s observations. The goal is to learn a policy $\pi_{\phi}(a\!\mid\!o)$ that maximizes the expected discounted return
@@ A.1 Partially-Observable Markov Decision Processes and Soft Actor-Critic
where $\mathcal{H}$ is the policy entropy and $\alpha$ a temperature hyperparameter. SAC alternates between minimizing the soft-Bellman residual for the Q-function,
@@
and updating the actor by minimizing the KL divergence to the soft Q-values.
@@ A.2 Contrastive Representation Learning
where $\mathrm{sim}(u,v)=u^{\top}v/\|u\|\|v\|$ and $\tau$ is a temperature. Once frozen, $f_{I}(x_{t})$ and a fixed goal embedding $f_{L}(y^{*})$ yield a dense, zero-shot reward:
@@
quantifying semantic alignment to a natural-language objective.
A.3 Chain-of-Thought Knowledge Distillation
@@ Appendix B Combined Reward and Training Algorithm
In this section we present the mathematical formulation of DriveMind’s composite reward, derive each component in detail, and lay out the complete training algorithm with all update rules. We begin by restating the per-step reward decomposition, then describe the Soft Actor-Critic updates, the world-model optimization, the chain-of-thought distillation for the dynamic VLM, and the novelty detection mechanism. Finally, we provide the complete pseudocode.
@@ B.1 Reward Decomposition
DriveMind’s per-step scalar reward $r_{t}$ combines four terms:
@@
  • $r_{\mathrm{task},t}$ is the environment-provided task reward (e.g. progress along route).
@@
  • $r^{\mathrm{AICR}}_{t}$ is the Adaptive ideal-State Contrastive Reward,
@@ B.2 Soft Actor-Critic Updates
We train a stochastic policy $\pi_{\phi}(a|o)$ and a soft Q-function $Q_{\theta}(o,a)$ via Soft Actor-Critic. The SAC objective consists of:
@@ B.3 World Model Optimization
The compact world model $M_{\phi}\colon\mathbb{R}^{k}\times\mathcal{A}\to\mathbb{R}^{k}$ is trained to minimize the one-step prediction error:
@@
We perform $U_{\mathrm{WM}}$ world-model updates for every SAC batch, using Adam with learning rate $\eta_{\mathrm{WM}}$.
B.4 Chain-of-Thought Distillation
When invoked, the dynamic VLM $\mathrm{VLM}^{D}$ is fine-tuned via CoT distillation from a GPT-4 teacher. Given an input image $x$, the teacher produces a target prompt sequence $y_{\rm prompt}$ and an accompanying chain-of-thought rationale $y_{\rm CoT}$. We update the student model by minimizing a combined cross-entropy loss over both the prompt and the reasoning tokens:
@@ B.5 Novelty Detection Formalism
The bird’s-eye-view (BEV) encoder $\mathrm{VLM}^{C}_{I}$ maps $\psi(s_{t})$ to $v_{t}\in\mathbb{R}^{k}$. We maintain a buffer =$\mathcal{B}=\{v^{(i)}\}_{i=1}^{K}$. The novelty score is:
@@ C.1.1 Reinforcement Learning Configuration
In our experiments, the DriveMind agent is trained with Soft Actor-Critic (SAC) as implemented in Stable-Baselines3. At each timestep $t$, the agent receives an observation
@@
  • $\psi(s_{t})\in\mathbb{R}^{H\times W\times C}$ is the bird’s-eye-view (BEV) semantic segmentation image of the environment.
  • $e_{t}=\bigl{(}v_{t},\;\theta_{t},\;\omega_{t}\bigr{)}\in\mathbb{R}^{3}$ is the ego-state vector, comprising the vehicle’s longitudinal speed $v_{t}$, steering angle $\theta_{t}$, and yaw rate $\omega_{t}$.
@@
[…]
@@
[…]
@@
  • BEV encoder $f_{\rm bev}\colon\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{d_{v}}$. A 6-layer convolutional network with ReLU activations:
@@
[…]
@@
  • Critic network $Q_{\theta}$: a three-layer MLP mapping $\bigl{(}z_{t},a_{t}\bigr{)}\to\mathbb{R}$.
  • Actor network $\pi_{\phi}$: a three-layer MLP mapping $z_{t}$ to the parameters of a diagonal Gaussian over $\mathcal{A}$.
@@
We optimize the soft Q-function by minimizing the Bellman residual:
@@
All networks are trained off-policy using a replay buffer $\mathcal{D}$, with target networks $Q_{\bar{\theta}}$ periodically updated.

### v2

# DriveMind: A Dual Visual Language Model-based Reinforcement Learning Framework for Autonomous Driving

@@ Abstract
[…]
@@ 1 Introduction
We introduce DriveMind, which comprises four key modules, each targeting a key limitation of existing approaches. First, a contrastive VLM provides dense, context-aware semantic rewards by anchoring each bird’s-eye-view (BEV) frame against fixed “present” and “ideal” concepts, addressing the lack of informative feedback and poor context adaptation in traditional reward designs. Second, a novelty-triggered VLM encoder–decoder is invoked only when embedding drift exceeds a threshold, generating adaptive “present” (hazard) and “ideal” (goal) prompts to prevent reward hacking in repetitive or adversarial scenarios. This on‐demand prompting with asynchronous chain-of-thought (CoT) distillation using GPT-4 minimizes per‐step computation, refreshing embeddings only upon genuine scene changes, hence having negligible impact on the overall latency while maintaining efficiency. Third, a hierarchical fusion module multiplicatively combines normalized kinematic factors, such as speed regulation, lane centering, heading alignment, lateral stability, to enforce a hard safety veto whenever any physical constraint is violated. Fourth, a compact world model forecasts the next semantic embedding and provides a predictive contrastive foresight reward, improving long-horizon credit assignment and anticipatory planning. Together, these modules yield an adaptive, interpretable, and provably safer reward framework for autonomous driving. Experiments on CARLA Town 2 demonstrate improved route completion, collision avoidance, and sample efficiency. Zero‐shot transfer to real dash-cam footage shows robust cross-domain alignment, confirming DriveMind’s promise for safer, more interpretable end‐to‐end driving.
@@
  • We validate DriveMind via extensive experiments on CARLA Town 2 and zero-shot transfer to BDD100K dash-cam data, achieving $19.4\pm 2.3$ km/h average speed, $0.98\pm 0.03$ route completion, near-zero collision speed in simulation, and minimal distributional shift in real data (Wasserstein =0.028=0.028, KS =0.105=0.105), surpassing baselines on most safety and efficiency metrics.
@@ 3 Proposed Approach: DriveMind
Figure 1: Dual‑VLM architecture and reward pipeline of DriveMind: A static contrastive VLM encoder maps each bird’s-eye-view image $\psi(s_{t})$ to a fixed semantic embedding $v_{t}$, while a novelty detector asynchronously triggers a dynamic VLM encoder-decoder to update “present” and “ideal” prompt embeddings $(u^{P}_{t},u^{I}_{t})$ via distilled GPT‑4 chain-of-thought. These embeddings feed into the Reward Combiner, yielding (i) Adaptive Ideal-State Contrastive Reward, (ii) Hierarchical Vehicle-State Fusion Reward, and (iii) Predictive Contrastive Foresight Reward.
@@ 3.4.3 Predictive Contrastive Foresight Module
Combining all components yields the final per‐step reward:
@@ 4.1 Experimental Setup
Observations & Actions. At each timestep, the agent receives an observation $o_{t}$ consisting of three parts. First, a bird’s-eye-view semantic image, resized to $224\times 224$ px, fuses camera and LiDAR segmentation of lanes, vehicles, pedestrians, and obstacles. Second, an ego-state vector $e_{t}=(v_{t},\theta_{t},\omega_{t})$ captures normalized longitudinal speed, steering angle, and yaw rate from onboard sensors. Third, a sequence of $K=10$ future waypoints $(x_{t+1},y_{t+1},\dots,x_{t+K},y_{t+K})$, sampled every 5 m in the vehicle’s local frame and scaled to [1,1][-1,1], provides route guidance. Together, these inputs form the full policy observation. The agent outputs a continuous action $a_{t}=(a^{(1)},a^{(2)})\in[-1,1]^{2}$, where $a^{(1)}$ maps linearly to steering angle and $a^{(2)}$ splits into throttle (positive) and brake (negative), each normalized to $[0,1]$. This compact interface enables real-time, smooth control in response to learned reward signals.
@@
[…]
@@ 4.3 Ablation Studies
DriveMind, NoHVFR $58\pm 0.25 0.0$$02\pm 0.00$ $14\pm 0.21 0.0$$00\pm 0.00$
DriveMind, NoPCFM $19.2\pm 1.14$ $92\pm 0.05 0.0$$02\pm 0.02$ $0.87\pm 0.06$
@@ 4.4 Real-World Generalization
Table 3 compares the Adaptive Ideal-State Contrastive Reward (AICR) distributions in CARLA Town 2 versus 10,000 BDD100K bird’s-eye-view frames. On real data, the mean reward drops slightly from 0.130 to 0.118 and the standard deviation rises modestly from 0.080 to 0.085. The fraction of positive‐reward frames falls from 88.0 % to 77.4 %, while negative‐reward frames increase from 12.0 % to 22.6 %. Crucially, the Wasserstein‐1 distance (0.028) and the Kolmogorov-Smirnov statistic (0.105) remain low, indicating that the two empirical distributions closely align. These results confirm that DriveMind’s learned semantic objectives transfer effectively, maintaining robust AICR behavior in unconstrained, real‐world driving scenes without any additional fine‐tuning.
@@
Dataset Mean $r_{\rm AICR}$ STD $r_{\rm AICR}$ % $r_{\rm AICR}>0$ % $r_{\rm AICR}<0$ EMD (CARLA $\rightarrow$ Real) KS-stat CARLA Town 2 0.130 0.080 88.0% 12.0% , , BDD100K (10K frames) 0.118 0.085 77.4% 22.6% 0.028 0.105
@@ Appendix A Preliminaries
A.1 Partially‐Observable Markov Decision Processes and Soft Actor‑Critic
We model autonomous driving as a partially‐observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},R,\gamma)$, where $\mathcal{S}$ is the hidden state space, $\mathcal{A}$ the continuous control actions, and $\mathcal{O}$ the agent’s observations. The goal is to learn a policy $\pi_{\phi}(a\!\mid\!o)$ that maximizes the expected discounted return
@@ A.1 Partially‐Observable Markov Decision Processes and Soft Actor‑Critic
where $\mathcal{H}$ is the policy entropy and $\alpha$ a temperature hyperparameter. SAC alternates between minimizing the soft‐Bellman residual for the Q‐function,
@@
and updating the actor by minimizing the KL divergence to the soft Q‐values.
@@ A.2 Contrastive Representation Learning
where $\mathrm{sim}(u,v)=u^{\top}v/\|u\|\|v\|$ and $\tau$ is a temperature. Once frozen, $f_{I}(x_{t})$ and a fixed goal embedding $f_{L}(y^{*})$ yield a dense, zero‐shot reward:
@@
quantifying semantic alignment to a natural‐language objective.
A.3 Chain‑of‑Thought Knowledge Distillation
@@ Appendix B Combined Reward and Training Algorithm
In this section we present the mathematical formulation of DriveMind’s composite reward, derive each component in detail, and lay out the complete training algorithm with all update rules. We begin by restating the per‐step reward decomposition, then describe the Soft Actor-Critic updates, the world‐model optimization, the chain‐of‐thought distillation for the dynamic VLM, and the novelty detection mechanism. Finally, we provide the complete pseudocode.
@@ B.1 Reward Decomposition
DriveMind’s per‐step scalar reward $r_{t}$ combines four terms:
@@
  • $r_{\mathrm{task},t}$ is the environment‐provided task reward (e.g. progress along route).
@@
  • $r^{\mathrm{AICR}}_{t}$ is the Adaptive ideal‐State Contrastive Reward,
@@ B.2 Soft Actor-Critic Updates
We train a stochastic policy $\pi_{\phi}(a|o)$ and a soft Q‐function $Q_{\theta}(o,a)$ via Soft Actor-Critic. The SAC objective consists of:
@@ B.3 World Model Optimization
The compact world model $M_{\phi}\colon\mathbb{R}^{k}\times\mathcal{A}\to\mathbb{R}^{k}$ is trained to minimize the one‐step prediction error:
@@
We perform $U_{\mathrm{WM}}$ world‐model updates for every SAC batch, using Adam with learning rate $\eta_{\mathrm{WM}}$.
B.4 Chain‐of‐Thought Distillation
When invoked, the dynamic VLM $\mathrm{VLM}^{D}$ is fine‐tuned via CoT distillation from a GPT-4 teacher. Given an input image $x$, the teacher produces a target prompt sequence $y_{\rm prompt}$ and an accompanying chain‐of‐thought rationale $y_{\rm CoT}$. We update the student model by minimizing a combined cross‐entropy loss over both the prompt and the reasoning tokens:
@@ B.5 Novelty Detection Formalism
The bird’s-eye-view (BEV) encoder $\mathrm{VLM}^{C}_{I}$ maps $\psi(s_{t})$ to $v_{t}\in\mathbb{R}^{k}$. We maintain a buffer $\mathcal{B}=\{v^{(i)}\}_{i=1}^{K}$. The novelty score is:
@@ C.1.1 Reinforcement Learning Configuration
In our experiments, the DriveMind agent is trained with Soft Actor-Critic (SAC) as implemented in Stable‑Baselines3. At each timestep $t$, the agent receives an observation
@@
  • $\psi(s_{t})\in\mathbb{R}^{H\times W\times C}$ is the bird’s‑eye‑view (BEV) semantic segmentation image of the environment.
  • $e_{t}=\bigl(v_{t},\;\theta_{t},\;\omega_{t}\bigr)\in\mathbb{R}^{3}$ is the ego‑state vector, comprising the vehicle’s longitudinal speed $v_{t}$, steering angle $\theta_{t}$, and yaw rate $\omega_{t}$.
@@
    Lateral deviation from the lane centerline exceeds $d_{\max}$ (e.g. $3\text{\,m}$).
@@
    Vehicle speed remains below $v_{\min}$ (e.g. $1\text{\,km/h}$) for more than $T_{\rm stuck}$ (e.g. $90$s).
@@
  • BEV encoder $f_{\rm bev}\colon\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{d_{v}}$. A 6‑layer convolutional network with ReLU activations:
@@
    Ego‐state encoder $f_{\rm ego}$. A two‐layer MLP that takes a 3‐dimensional ego state vector $(\text{steering},\;\text{throttle},\;\text{speed})$ as input and outputs a $d_{e}$‐dimensional embedding:
@@
  • Critic network $Q_{\theta}$: a three‑layer MLP mapping $\bigl(z_{t},a_{t}\bigr)\to\mathbb{R}$.
  • Actor network $\pi_{\phi}$: a three‑layer MLP mapping $z_{t}$ to the parameters of a diagonal Gaussian over $\mathcal{A}$.
@@
We optimize the soft Q‑function by minimizing the Bellman residual:
@@
All networks are trained off‑policy using a replay buffer $\mathcal{D}$, with target networks $Q_{\bar{\theta}}$ periodically updated.